OpenStudy (anonymous):

2 MnO(s) + 5 PbO2(s) + 10 HNO3(aq) 2 HMnO4(aq) + 5 Pb(NO3)2(aq) + 4 H2O(l) howmany moles of water would be produced if 13.2 mol of lead (II) nitrate were formed

OpenStudy (aaronq):

Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\)